# Learning from Label Proportions: Bootstrapping Supervised Learners via Belief Propagation

### 17 Jan 2024 v3

@@ 1 Introduction
Learning from Label Proportions (henceforth LLP) has seen renewed interest in recent times due to the rising concerns of privacy and leakage of sensitive information (Ardehaly & Culotta, 2017; Busa-Fekete et al., 2023; Zhang et al., 2022; Kobayashi et al., 2022; Yu et al., 2014). In the LLP binary classification setting, all the training instances are aggregated into bags and only the aggregated label count for a bag is available, i.e. proportion of $1$’s in a bag. Features of all instances are available. This can be seen as a form of weak supervision compared to providing instance-level labels. The main goal is learn an instance wise predictor that performs very well on the test distribution.
@@
Our work builds on the idea of forming pseudo labels per instance. Our key observation is that one could utilize two types of information: bag level constraints and covariate similarity information in an explicit way. To realize this, we take inspiration from coding theory for communication systems (MacKay & Mac Kay, 2003; Kschischang et al., 2001), where one of the fundamental problems is to decode an unknown message string sent by the encoder using only parity checks over groups of bits from the message. Parity checks provide redundancy that battles against noise/corruptions in the channel. The state of the art codes (MacKay & Mac Kay, 2003) are decoded using Belief Propagation (Pearl, 1988) on a factor graph where message bits are variable nodes (whose "label" is to be learnt) and parity checks form factor nodes (that constrain the sum of bits in the parity checks to be odd or even). Sum-product belief propagation is used to learn the marginal soft score on the each of the message bits.
@@ 3 Problem Setting and Overview of Our Solution
Consider a supervised learning dataset ${\cal D}=\{(\mathbf{x}_{i},y_{i})_{i=1}^{m}\}$ where $x_{i}\in\mathbb{R}^{d},~{}y_{i}\in\{0,1\}$. The instance wise labels $y_{i}$’s are not explicitly revealed to the learning agent. There are a set of "bags" =${\cal B}=\{S_{1}\ldots S_{n}\}$ that contains subsets $S_{i}\subseteq[m]$ and for each $S_{i}$, bag level counts $y(S_{i})=\sum_{j\in S_{i}}y_{j}$. In this work, we will consider the case of disjoint bags, i.e. $S_{i}\cap S_{j}=\emptyset$. Let the vector of bag level counts be $\left[y(S_{1})\ldots y(S_{n})\right]=y({\cal B})$. In this work, we consider the following problem:
@@ 4.1 Step 1: Obtaining Pseudo-Labels through Belief Propagation (BP)
Remark: We note that not all pairwise terms are present. If two points are not in K-NN neighborhood of each other and if they don’t belong together in any bag, then there would be no pairwise term corresponding to it. In our experiments, we consider the case where Bags are disjoint and non overlapping. Therefore, every instance $i$ participates in only one bag. Use of K-NN and small disjoint bags creates only linear number of terms in the Gibbs distribution.
@@
Obtaining Pseudo Labels using sum-product Belief Propagation(BP): We use the classical sum-product Belief Propagation (MacKay & Mac Kay, 2003) on equation 4.1 to approximate the marginal distribution $\mathbb{P}_{\lambda_{s},\lambda_{b}}(y_{i})$. We briefly describe the algorithm, At every round $t$, node $i$ passes the following message $m^{t}_{j\rightarrow i}(y_{i}),~{}y_{i}\in\{0,1\}$ to every node $j:J_{i,j}\neq 0$ given by:
@@ 5.1 Experimental Setup
We optimize our algorithm, using hyperparameters $\lambda_{s},\lambda_{b}\in[10^{-4},200]$, $k\in[1,30]$, $T=[50,100,200]$, $\tau\in(0,1),MLP_{LR}\in[10^{-6},1],MLP_{WD}\in[10^{-12},10^{-1}],\lambda_{a}\in[0,10],\delta_{d}\in[10^{-4},1],BatchSize_{train}=[2,4,8,\ldots 4096,8192]$, tuned using Vizier (Song et al., 2022) to achieve the best Validation AUC score. Illustrative values of best hyperparameters for various experiments are given in appendix section A.6. We then report the corresponding Test AUROC % averaged over 3 trials and report the sample standard deviation in parenthesis throughout our tables. We perform the same setup for all our baselines. Best number is reported in bold and 2nd best is reported in underline
@@ 5.2 Performance Analysis
Our method scales really well for $1$ million samples with negligible computational cost for the sum-product BP step. Criteo is an inherently harder dataset to work with due to high feature dimensionality and most of them being categorical. Over the last 7 years, the dataset has seen a 2% improvement in AUC while our method is able to produce a 0.8% improvement over DLLP. Similar to the previous tabular datasets, we also observe here that the iteration seem to help improve performance.
@@
Among the 2 image datasets, CIFAR-S has a large label skew. Our methods outperforms all baseline for CIFAR-S in the large bag size regime ($BagSize\geq 512$), where the performance of all other methods drop significantly. Specifically, we outperform DLLP by up to 20% and GenBags by up to 7%. For small bags, our method performs comparable to the best baseline.
Only for CIFAR-B that has label balance, GenBags is better than our method for larger bag sizes while DLLP outperforms slightly for lower bag sizes. However, even in this case, our method is competitive (close second mostly) with the best across bag sizes.
@@ 6 Ablations
Here we provide ablations regarding the two most important ideas in our algorithm: 1) Nearest neighbor based nearness constraints and 2) Aggregate Embedding. We do provide a number of other ablations in the appendix regarding adding noise to embeddings (A.1.2), hard vs soft thresholding of BP labels (A.1.3), performance change with different choice of kernels (A.1.5) and distance functions (A.1.4) among others. In the supplement, we further report various performance metrics of our algorithm such as performance of BP pseudo labels in itself compared to ground truth (Section A.2.3) and convergence in very few iterations (typically 1-2) (Section A.2.2).
@@ 6.1 Time Complexity
Table 4: Time for various parts of our algorithm compared to time taken by DLLP on Criteo Dataset. All time values are in seconds.
@@
Bag Size DLLP Training Ours - Data Setup Ours - BP Ours - MLP Ours - Total
8 725.24 (142.59) 1993.47 (96.05) 695.34 (266.7) 679.26 (215.35) 3368.07 (466.34)
32 735.09 (207.50) 1970.05 (104.65) 1279.83 (278.75) 624.55 (187.30) 3874.43 (469.92)
128 568.00 (79.14) 2192.84 (189.90) 3590.73 (528.85) 588.00 (157.84) 6371.57 (727.31)
@@ 7 Conclusion
We have described in detail the implementation details for the reproducibility of the experiments in the main paper Section 5.1 supplementary material Section A.5. We provide the hyperparameter ranges and experimental methodology in section 5.1 and additional information, including that for the baselines in supplementary section A.5. We have provided extensive set of hyperparameters in A.6 to reproduce our algorithm’s numbers. We will publicly release the source code soon.
@@ A.2.1 Wall Clock Times on Adult Dataset
Table 5: Time for various parts of our algorithm compared to time taken by DLLP on Adult Dataset. All time values are in seconds.
@@
Bag Size DLLP Training Ours - Data Setup Ours - BP Ours - MLP Ours - Total
8 52.23 (51.38) 630.88 (827.63) 12.34 (8.57) 85.00 (145.57) 728.21 (863.69)
32 45.4 (69.04) 523.59 (449.32) 21.11 (14.65) 54.38 (77.15) 599.08 (458.12)
128 73.79 (125.96) 623.20 (458.23) 71.16 (8.89) 54.34 (63.55) 748.71 (459.86)
512 25.11 (40.15) 767.58 (431.50) 269.88 (32.52) 65.07 (158.34) 1102.53 (459.12)
1024 26.39 (38.06) 645.04 (674.94) 550.32 (113.28) 32.42 (25.76) 1227.78 (756.34)
2048 24.34 (28.32) 612.87 (539.07) 1054.47 (73.71) 49.93 (78.34) 1717.27 (534.88)
@@ A.2.4 Noisy Labels and Privacy
Note: We will update the citations in the main paper to include the updated citations mentioned as footnote in the supplementary section.
@@ Appendix B Extended Related Work
Belief Propagation: Belief Propagation (BP) has been used to compute marginals and find MAP estimates in standard sparse graphical models, like Bayesian networks and Markov random fields (Pearl, 2022) by message passing across edges on an appropriate graph. Sum-product BP algorithm is used for computing marginals and it is known to converge on trees. It was also extended to polytrees (Kim & Pearl, 1983). It is also an effective approximate algorithm on general graphical models (Pearl, 1988). More relevant to our work is the fact that sum product Belief Propagation has found widespread in communication system, where it is used to soft-decode a binary string message from their parity checks as in LDPC (low-density parity-check) codes (Richardson & Urbanke, 2001; Gallager, 1962), iterative decoding of turbo codes (MacKay & Mac Kay, 2003; Kschischang et al., 2001). In communication codes, parity checks are designed so as to have nice properties on the graphical models they induce. In our problem, the bag levels constraints can be thought of as parity checks but are given and we add additional constraints from covariate information that is also given. We use a public scalabale and efficient implementation PGMax (Zhou et al., 2022) of the sum-product message passing algorithm.
@@
Appendix C Limitations and Future Work

### 20 Mar 2024 v4

@@ Abstract
00footnotetext: $*{}^{*}$qual Contribution. Correspondence to {shreyasjh,navoditasharma}@google.com
@@ 1 Introduction
Learning from Label Proportions (henceforth LLP) has seen renewed interest in recent times due to the rising concerns of privacy and leakage of sensitive information (Ardehaly & Culotta, 2017; Busa-Fekete et al., 2023; Zhang et al., 2022; Kobayashi et al., 2022; Yu et al., 2014; Chen et al., 2023). In the LLP binary classification setting, all the training instances are aggregated into bags and only the aggregated label count for a bag is available, i.e. proportion of $1$’s in a bag. Features of all instances are available. This can be seen as a form of weak supervision compared to providing instance-level labels. The main goal is learn an instance wise predictor that performs very well on the test distribution.
@@
Our work builds on the idea of forming pseudo labels per instance. Our key observation is that one could utilize two types of information: bag level constraints and covariate similarity information in an explicit way. To realize this, we take inspiration from coding theory for communication systems (MacKay, 2003; Kschischang et al., 2001), where one of the fundamental problems is to decode an unknown message string sent by the encoder using only parity checks over groups of bits from the message. Parity checks provide redundancy that battles against noise/corruptions in the channel. The state of the art codes (MacKay, 2003) are decoded using Belief Propagation (Pearl, 1988) on a factor graph where message bits are variable nodes (whose "label" is to be learnt) and parity checks form factor nodes (that constrain the sum of bits in the parity checks to be odd or even). Sum-product belief propagation is used to learn the marginal soft score on the each of the message bits.
@@ 3 Problem Setting and Overview of Our Solution
Consider a supervised learning dataset ${\cal D}=\{(\mathbf{x}_{i},y_{i})_{i=1}^{m}\}$ where $x_{i}\in\mathbb{R}^{d},~{}y_{i}\in\{0,1\}$. The instance wise labels $y_{i}$’s are not explicitly revealed to the learning agent. There are a set of "bags" =${\cal B}=\{S_{1}\ldots S_{n}\}$ that contains subsets $S_{i}\subseteq{[m]}$ (thus denoting the indices that correspond to the instances present in that "bag") and for each $S_{i}$, bag level counts $y(S_{i})=\sum_{j\in S_{i}}{y^{true}_{j}}$. In this work, we will consider the case of disjoint bags, i.e. $S_{i}\cap S_{j}=\emptyset$. Let the vector of bag level counts be $\left[y(S_{1})\ldots y(S_{n})\right]=y({\cal B})$. In this work, we consider the following problem:
@@ 4.1 Step 1: Obtaining Pseudo-Labels through Belief Propagation (BP)
[…]
@@
Obtaining Pseudo Labels using sum-product Belief Propagation (BP): We use the classical sum-product Belief Propagation (MacKay, 2003) on equation 4.1 to approximate the marginal distribution $\mathbb{P}_{\lambda_{s},\lambda_{b}}(y_{i})$. We briefly describe the algorithm, At every round $t$, node $i$ passes the following message $m^{t}_{j\rightarrow i}(y_{i}),~{}y_{i}\in\{0,1\}$ to every node $j:J_{i,j}\neq 0$ given by:
@@ 5.1 Experimental Setup
We optimize our algorithm, using hyperparameters $\lambda_{s},\lambda_{b}\in[10^{-4},200]$, $k\in[1,30]$, $T=[50,100,200]$, $\tau\in(0,1),MLP_{LR}\in[10^{-6},1],MLP_{WD}\in[10^{-12},10^{-1}],\lambda_{a}\in[0,10],\delta_{d}\in[10^{-4},1],BatchSize_{train}=[2,4,8,\ldots 4096,8192]$, tuned using Vizier (Song et al., 2022) to achieve the best Validation AUC score. Illustrative values of best hyperparameters for various experiments are given in appendix section A.6. We then report the corresponding Test AUROC % averaged over 3 trials and report the sample standard deviation in parenthesis throughout our tables. We perform the same setup for all our baselines. Best number is reported in bold and 2nd best is reported in underline.
@@ 5.2 Performance Analysis
Our method scales really well for $1$ million samples with negligible computational cost for the sum-product BP step. Criteo is an inherently harder dataset to work with due to high feature dimensionality and most of them being categorical. Over the last 7 years, the dataset has seen a 2% improvement in AUC while our method is able to produce a 0.8% improvement over DLLP. Similar to the previous tabular datasets, we also observe here that the iteration seems to help improve performance.
@@
Among the 2 image datasets, CIFAR-S has a large label skew. Our methods outperforms all baseline for CIFAR-S in the large bag size regime ($BagSize\geq 512$), where the performance of all other methods drop significantly. Specifically, we outperform DLLP by upto 20% and GenBags by upto 7%. For small bags, our method performs comparable to the best baseline. Only for CIFAR-B that has label balance, GenBags is better than our method for larger bag sizes while DLLP outperforms slightly for lower bag sizes. However, even in this case, our method is competitive (close second mostly) with the best across bag sizes.
Note: We justify our chosen hyper-parameters in our experiments via approximate convergence analysis in the supplementary section C providing theoretical backing for our strong empirical results.
@@ 6 Ablations
Here we provide ablations regarding the two most important ideas in our algorithm: 1) Nearest neighbor based nearness constraints and 2) Aggregate Embedding. Due to space constraints, we do provide a number of other ablations in the appendix regarding adding noise to embeddings (A.1.2), hard vs soft thresholding of BP labels (A.1.3), performance change with different choice of kernels (A.1.5) and distance functions (A.1.4) among others. In the supplement, we further report various performance metrics of our algorithm such as performance of BP pseudo labels in itself compared to ground truth (Section A.2.3) and convergence in very few iterations (typically 1-2) (Section A.2.2).
@@ 6.1 Time Complexity
Table 4: Time for various parts of our algorithm compared to time taken by other methods on Criteo Dataset. All time values are in seconds. Note that Data Setup time is common to All Methods
@@
Bag Size DLLP Training EasyLLP Training GenBags Training Data Setup Ours - BP Ours - MLP Ours - Total
8 725.24 (142.59) 1617.67 (1597.14) 1760.33 (1077.05) 1993.47 (96.05) 695.34 (266.7) 679.26 (215.35) 3368.07 (466.34)
32 735.09 (207.50) 911.86 (553.21) 957.65 (497.01) 1970.05 (104.65) 1279.83 (278.75) 624.55 (187.30) 3874.43 (469.92)
128 568.00 (79.14) 777.82 (530.75) 729.41 (502.12) 2192.84 (189.90) 3590.73 (528.85) 588.00 (157.84) 6371.57 (727.31)
@@ 7 Conclusion
We have described in detail the implementation details for the reproducibility of the experiments in the main paper Section 5.1 supplementary material Section A.5. We provide the hyperparameter ranges and experimental methodology in section 5.1 and additional information, including that for the baselines in supplementary section A.5. We have provided extensive set of hyperparameters in A.6 to reproduce our algorithm’s numbers. We will soon publicly release the source code.
@@ A.2.1 Wall Clock Times on Adult Dataset
Table 5: Time for various parts of our algorithm compared to time taken by other methods on Adult Dataset. All time values are in seconds. Note that Data Setup time is common to All Methods
@@
Bag Size DLLP Training EasyLLP Training GenBags Training Data Setup Ours - BP Ours - MLP Ours - Total
8 52.23 (51.38) 47.85 (28.93) 256.94 (495.04) 630.88 (827.63) 12.34 (8.57) 85.00 (145.57) 728.21 (863.69)
32 45.4 (69.04) 26.51 (12.67) 125.67 (102.96) 523.59 (449.32) 21.11 (14.65) 54.38 (77.15) 599.08 (458.12)
128 73.79 (125.96) 21.38 (15.99) 88.72 (76.21) 623.20 (458.23) 71.16 (8.89) 54.34 (63.55) 748.71 (459.86)
512 25.11 (40.15) 18.84 (9.42) 85.52 (74.49) 767.58 (431.50) 269.88 (32.52) 65.07 (158.34) 1102.53 (459.12)
1024 26.39 (38.06) 20.93 (14.51) 95.79 (103.85) 645.04 (674.94) 550.32 (113.28) 32.42 (25.76) 1227.78 (756.34)
2048 24.34 (28.32) 18.77 (6.69) 86.14 (71.80) 612.87 (539.07) 1054.47 (73.71) 49.93 (78.34) 1717.27 (534.88)
@@ A.2.4 Noisy Labels and Privacy
A.2.5 MAP Decoding
Table 9 and Table 10 denote the Test AUC % on performing Max Product BP on the Gibbs Distribution, and highlight the noisy nature of the performance. This establishes the superiority of Sum Product BP, that is our approach, to obtain consistently good performance across bag sizes and datasets. On the other hand, the single label configuration obtained from the Max Product approach is noisy and does not consistently retrieve the same performance as in the Sum Product approach across bag sizes.
One plausible reason is that given the nature of the weak supervision, i.e. aggregate labels being available only at the level of bags, it is better to find out uncertainty in a label for an instance (marginalize) rather than commit to a MAP configuration with respect to a Gibbs distribution that is uncertain.
Table 9: MaxProduct BP on Adult Dataset
Bag Size Test AUC Itr-1 Test AUC Itr-2
8 89.15 88.92
32 75.29 87.9
128 76.08 84.67
512 73.77 73.82
1024 80.43 76.65
2048 74.45 68.69
Table 10: MaxProduct BP on Marketing Dataset
Bag Size Test AUC Iter 1 Test AUC Iter 2
8 85.55 85.73
32 83.99 83.67
128 82.64 82.39
512 72.29 80.66
1024 72.48 73.03
2048 72.49 80.42
A.2.6 Extension to the Multiclass Paradigm
We adapted the Gibbs measure to the multi class setting as follows. Every point has $k$ labels : $y^{1}_{i}\ldots y^{k}_{i}$ corresponding to $k$ classes.
We have three main types of terms in the Gibbs measure. We impose a soft one hot constraint with the term: $(\sum_{p}y^{p}_{i}-1)^{2}$
Nearness terms get modified as follows: $K(x_{i},x_{j})\sum_{p}(y^{p}_{i}-y^{p}_{j})^{2}$, i.e. Euclidean distance between the vector labels of two points is small if they are nearby.
Aggregate Bag level constraints have counts $b_{1}\ldots b_{k}$. Then we simply impose a least squares constraint: $\sum_{p}(\sum_{i\in B}y^{p}_{i}-b_{p})^{2}$
All these terms are scaled by temperature hyper parameters which we search over during training. Essentially these are vectorized least squares constraints.
The results on CIFAR10 as provided in Table 11 depict that our method is slightly better or comparable to the SOTA.
Table 11: % Accuracy of our method against high performing baselines on CIFAR10 multiclass classification.
Bag Sizes 8 32
LLP-VAT 66.91 (1.23) 60.85 (2.45)
LLP-FC 66.52 (2.12) 62.35 (1.32)
DLLP 68.14 (0.48) 62.87 (0.88)
Ours-Itr-1 68.05 (0.23) 61.83 (0.21)
Ours-Itr-2 69.23 (0.12) 61.92 (0.43)
@@ Appendix B Extended Related Work
Belief Propagation: Belief Propagation (BP) has been used to compute marginals and find MAP estimates in standard sparse graphical models, like Bayesian networks and Markov random fields (Pearl, 2022) by message passing across edges on an appropriate graph. Sum-product BP algorithm is used for computing marginals and it is known to converge on trees. It was also extended to polytrees (Kim & Pearl, 1983). It is also an effective approximate algorithm on general graphical models (Pearl, 1988). More relevant to our work is the fact that sum product Belief Propagation has found widespread in communication system, where it is used to soft-decode a binary string message from their parity checks as in LDPC (low-density parity-check) codes (Richardson & Urbanke, 2001; Gallager, 1962), iterative decoding of turbo codes (MacKay, 2003; Kschischang et al., 2001). In communication codes, parity checks are designed so as to have nice properties on the graphical models they induce. In our problem, the bag levels constraints can be thought of as parity checks but are given and we add additional constraints from covariate information that is also given. We use a public scalabale and efficient implementation PGMax (Zhou et al., 2022) of the sum-product message passing algorithm.
@@
Weak Supervision: There has been recent interest in exploring training of models under weak supervision, where complete label information is not available. Though these methods do not exactly map to the learning from label proportions setup, the idea of using pseudo labels to train a model with pre-trained representations has been explored before in Chen et al. (2022), Pukdee et al. (2022), Ratner et al. (2017), Karamanolakis et al. (2021), which creatively bypass the lack of labels, by either using covariate information, label propagation based on the few available labels, heuristically creating new labels or even using student-teacher models. While such approaches are designed to work in the case of lack of all instance labels, none of them deal with the specific case of weak supervision we concern ourselves with, namely learning only from the aggregate bag labels and no instance labels whatsoever. This is what makes the LLP setup even harder and sparser in terms of information available for the learner.
Appendix C Approximate Convergence Analysis
Loopy Belief Propagation in the literature: We would like to point out that even classical literature from the past Frey & MacKay (1997) has pointed out that while loopy belief propagation in graphs with cycles don’t have known convergence guarantees, in many applications like error correction for communicating over noisy channels, belief propagation based decoding perform extremely well (abstract of Frey & MacKay (1997) makes exactly this case). Even considering very recent work on message passing on complex networks Newman (2023), the conclusion remains that loopy belief propagation is effective in practice but not very easily amenable to analysis. Although we showed the edges due to the similarity constraints form a cycle free graph for $1$-NN based Gibbs distribution, bag constraints introduce cycles. However, Newman (2023) offers an approximate analysis based on linearized version of sum-product-BP that we adopt and we show that the inverse temperature parameters chosen for Adult Dataset (Table 12) are roughly orderwise within the stability region for an approximate linearized version of BP with the same graph structure as in the Gibbs distribution we define.
Sufficient conditions for convergence of the Sum-Product Algorithm (Mooij & Kappen, 2007):
We refer to Corollary 1 from (Mooij & Kappen, 2007) for a sufficient condition such that message updates from Loopy BP is a contraction. We substitute the values of $J_{ij}$, from our Gibbs Distribution to evaluate this condition which is: $\max\limits_{i}[(|N(i)|-1)\max\limits_{j\in N(i)}\tanh(|J_{ij}|)]<1$. For the case of using $1$-NN constraints, $|N(i)|-1=B-1$ ($B$ is the bag size). Now note that $\lvert J_{ij}\rvert\leq 2\lambda_{b}+4\lambda_{s}$ as discussed above and as observed from Table 12.
And we observe that for entries in Table 12 for example for bag size 8, for Corollary 1, the condition (Eq. 15 from the paper) is satisfied and thus our LBP is a $l_{\infty}$-contraction and converges to a unique fixed point, irrespective of the initial messages.
For larger bags, we offer an approximate linearized BP analysis below.
Stability analysis of the approximate linearized BP: Now, we offer some approximate analysis on convergence of the Belief Propagation step for some simpler cases (like the Adult Dataset). In what follows, we follow the recipe given in Newman (2023) to linearize the BP. We have a Gibbs distribution over $n$ binary variables $\mathbf{y}\in\{-1,+1\}^{n}$ given by:
$\mathbb{P}(\mathbf{y})\propto\exp\left(\sum\limits_{i}h_{i}y_{i}+\sum\limits_{i\neq j}J_{ij}y_{i}y_{j}\right)$
$h_{i}$ forms the external field that biases the variables away from $1/2$. However, we will analyze the Gibbs distribution without $h_{i}$ as this external field can be taken to be the prior bias for each of the variables.
$\{J_{ij}\}$ is very sparse and is non zero only if $i,j$ are in each others 1-NN neighborhood or $i,j$ belong to the same bag. $J_{ij}$ in our formulation could be negative but we actually, analyze the ferromagnetic version with only the correct adjacency matrix, i.e.
$\mathbb{P}(\mathbf{\tilde{y}})\propto\exp(-\beta\sum\limits_{i\neq j}\mathbf{1}_{J_{ij}\neq 0}\tilde{y}_{i}\tilde{y}_{j})$
For the above Ising model over $\mathbf{\tilde{y}}$, we have the following normalized update rule (this is from Newman (2023)):
$\frac{1}{2}(1+\epsilon_{i\leftarrow j})$ $=\frac{1}{Z_{i\leftarrow j}}\prod\limits_{k\in N(j)-i}\frac{1}{2}\left[\exp(\beta)(1+\epsilon_{j\leftarrow k})+\exp(-\beta)(1-\epsilon_{j\leftarrow k})\right]$
$Z_{i\leftarrow j}$ $=\sum\limits_{r=\{+1,-1\}}\prod\limits_{k\in N(j)-i}\frac{1}{2}\left[\exp(\beta r)(1+\epsilon_{j\leftarrow k})+\exp(-\beta r)(1-\epsilon_{j\leftarrow k})\right]$ (10)
Here, $N(j)$ is the neighborhood of $j$ according to the graph obtained from the $1$-00 adjacency matrix $1_{J_{i,j}\neq 0}$. This is equivalent to the BP iterations quoted in Section 4.1 (except for the normalization $Z_{i\leftarrow j}$ and setting $m_{i\leftarrow j}(+1)=1+\epsilon_{i\leftarrow j},~{}m_{i\leftarrow j}(-1)=1-\epsilon_{i\leftarrow j}$ due to the normalization.
Again following Newman (2023), and ignoring the second order terms in $O(\epsilon_{\{\cdot\}}^{2})$, we have the following linearized Belief Propagation:
$\epsilon_{i\leftarrow j}=\tanh{\beta}\sum\limits_{k\in N(j)-i}\epsilon_{j\leftarrow k}$ (11)
We have an edge - incidence matrix $H$ of size $2|E|\times 2|E|$ where $|E|$ is the set of edges in the graph indexed by valid oriented edges $i\leftarrow j$. For the row $i\leftarrow j$, all columns corresponding to $j\leftarrow k:k\neq i,~{}k\in N(j)$ have $1$ and all other entries are 00.
Essentially, the messages if the linearized BP converge, then it is a fixed point of the equation: $X=\tanh(\beta)Hx$. Therefore, convergence is exponentially fast if $\tanh(\beta)\lVert H\rVert_{2}<1$ as the linear operator becomes a contraction. Here, $\lVert H\rVert_{2}$ is the spectral norm of $H$.
Computation of inverse temperature threshold for Adult Dataset: For the full Adult dataset, the edge incidence matrix even for $1$-NN graph is of the order of $100$k. Hence, computing the largest singular value of this is a time consuming operation. However, we randomly subsampled $5k$ points from the distribution and we analyze the spectral norm of the edge incidence matrix obtained from this.
We found that $\lVert H\rVert_{2}\approx 6.344$. This means that inverse temperature is at most $\tanh^{-1}(1/6.344)\approx 0.158$ for the linearized BP to converge. We observe from Table 12 that $\lvert J_{ij}\rvert\leq 2\lambda_{b}+4\lambda_{s}$ is much smaller than this threshold for larger bags.
Key Takeaway: Exact loopy BP updates are shown to be a contraction based on sufficient conditions in Mooij & Kappen (2007) for smaller bag sizes for Adult Dataset. With an approximate linearized BP analysis of the ferromagnetic model with the same graph structure, for a subsampled set of data points i.i.d from the original Adult Dataset without replacement, we show that for large bags the chosen hyperparameters for inverse temperatures are well within the convergence threshold.
Appendix D Limitations and Future Work